[Brahms-l] Shift report 20040112 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 12 2004 - 00:07:29 EST
------------------------------------------------------------
Date: Mon Jan 12 00:07:28 2004, Shift: 16:00-24:00
Supervisor: Catalin Ristea, with: 
Log book#15 pages 30 - 30

Shift Summary:
------------------------------------------------------------
1. Good beam from 1800 until now.
2. We got the runs 9555, 9556, 9557, 9558 which give us a total of 1,000k events on trigger 6, with FS at 4degs and the magnetic field at 1/sqrt(2), A pol. Also took pedestal runs for all detectors.
3. Everything worked suspiciously well during this shift.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 12 00:07:45 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 12 2004 - 00:08:03 EST